# Joint learning of a network of linear dynamical systems via total variation penalization [^1] Authors are listed alphabetically

### v1

@@ 2.3 Graph–geometry parameters
the combinatorial Laplacian and let $L^{\dagger}$ be the pseudoinverse of $L$. Denote the eigenvalues of $\lambda_{1}(G)\geq\lambda_{2}(G)\geq\cdots\geq\lambda_{m-1}(G)>\lambda_{m}(G)(=0)$. Throughout, $\lambda_{m-1}(G)$ (or simply $\lambda_{m-1}$) stands for the algebraic connectivity—the smallest positive eigenvalue of $L$ (also known as the Fiedler eigenvalue of $G$). It is well known that $\lambda_{m-1}>0$ iff $G$ is connected, and more generally, $\lambda_{m-1}$ measures how well-knit the graph is. A small value indicates bottlenecks or sparse regions, making global information propagation harder.
@@ 2.4 Main results
which then allows us to suitably bound the terms $F_{1},F_{2}$ and $F_{3}$; see Section B for details.
@@
We now have the following useful corollary of Theorem 1 when each $A^{*}_{l}$ is stable. Its proof is outlined in Section B.
@@
    $\frac{1}{\sqrt{m}}\|\hat{a}-a^{*}\|_{2}\;\lesssim\;\frac{\mathfrak{M}}{\sqrt{mT}}\Big(1+\frac{\sqrt{s}}{\kappa_{S}}\Big),\qquad\Phi_{S}=1+\frac{\sqrt{s}}{\kappa_{\mathcal{S}}}\;.$
@@
$T\gtrsim\max\left\{{\frac{\log^{2}(1/\delta)}{\Delta^{2}}\Phi_{S}^{2}\log(dm/\delta),\frac{d^{2}+\log^{4}(1/\delta)}{m\,\Delta^{2}},\frac{\Phi_{S}\,d\,\Delta_{G}}{\sqrt{m}}}\right\},$
where we recall $\Phi_{S}=1+\|\widetilde{D}\,a^{*}\|_{1}$.
@@
$T\gtrsim\max\left\{{\frac{\log^{2}(1/\delta)}{\Delta^{2}}\Phi_{S}^{2}\log^{2}(dm/\delta),\frac{d^{2}+\log^{4}(1/\delta)}{m\,\Delta^{2}},\Phi_{S}\,d\,\Delta_{G}\sqrt{\frac{\log m}{m}}}\right\},$
where $\Phi_{S}=1+\|\widetilde{D}\,a^{*}\|_{1}$. As before, consider for simplicity the very smooth regime where $\Phi_{S},\Delta_{G}\leq c$ for a constant $c\geq 1$. Then if $\frac{m}{\log m}\gtrsim d^{2}+\log^{4}(1/\delta)$, we require $T\gtrsim\frac{\log^{2}(1/\delta)}{\Delta^{2}}\log^{2}(dm/\delta)$.
@@ 4 Experiments
In this section, we validate the performance of the graph total variation regularizer for estimating the LDS matrices $A^{*}_{l}$ on a series of synthetic experiments. To this end, we consider linear dynamical systems on a graph on $m$ nodes where each node $l\in\{1,\dots,m\}$ is associated with a time series that evolves as
@@ 4.1 Experimental Protocol
We vectorize each $A^{*}_{l}\in\mathbb{R}^{d\times d}$ and stack them in a single vector $a^{*}=\big(\operatorname{vec}(A^{*}_{1})^{\top},\dots,\operatorname{vec}(A^{*}_{m})^{\top}\big)^{\top}\in\mathbb{R}^{md^{2}}$. For node $l$, with time series $X_{l}\in\mathbb{R}^{d\times T_{\text{train}}}$ and response $\widetilde{X}_{l}\in\mathbb{R}^{d\times T_{\text{train}}}$, the local design is $Q_{l}=(X_{l}^{\top}\otimes I_{d})\in\mathbb{R}^{(T_{\text{train}}d)\times d^{2}}$. We form $Q=\operatorname{blkdiag}(Q_{1},\dots,Q_{m})$ and $\tilde{x}=\operatorname{vec}(\widetilde{X}_{l})\in\mathbb{R}^{dT_{\text{train}}}.$
@@
Given a graph and a collection of LDS matrices $(A^{*}_{l})_{l=1}^{m}$, we simulate $T_{\text{tot}}=T_{\text{train}}+2b+T_{\text{val}}+T_{\text{test}}$ steps per node from (4.1), with $b$ as a buffer parameter (here, chosen to be 100) that ensures that the validation and testing sets are (approximately) independent of one another and of the training set. We use the first $T_{\text{train}}$ steps for training, $T_{\text{val}}$ steps for validation, and the last $T_{\text{test}}$ for testing. Unless stated otherwise, we set $T_{\text{val}}=4$, $T_{\text{test}}=8$, and repeat over seeds $r=1,\dots,n_{\text{rep}}$.
@@ 4.2 Results
We also vary $T_{\text{train}}\in\{4,8,16,32,64,128,256\}$ (log-scale $x$-axis). We expect TV to clearly outperform OLSind at small $T$ by borrowing strength across nodes, while approaching OLSind as $T_{\text{train}}$ grows; OLSpooled is expected to perform well when $T_{\text{train}}$ is extremely small, as the benefits of the smoothing induced by fitting the estimator over the entire graph are expected to outweigh the bias incurred in this case. The results for the different topologies and signals are reported in Figures 5(a) and 5(b), and confirm our expected findings. In particular, for the path graph and dense Erdős–Rényi graph, the TV regularizer consistently outperforms all other estimators, including the Laplacian estimator. We also note that for the piecewise-constant signal, the graph TV estimator offers significant advantages in terms of estimation MSE over the Laplacian estimator (e.g. in the path and ER cases). Even in the 2D grid, as $T_{\text{train}}$ increases, we see that the TV regularizer improves upon the Laplacian. This is expected, as the TV regularizer’s ability to preserve sharp contrasts is advantageous for signals defined by a few sparse, high-magnitude differences rather than small, distributed ones.
(a) Parameter and test prediction MSE as a function of training time length $T$ and graph topology. The signal generated in these plots is piecewise constant (as per Equation 4.3), with $\|D\beta^{*}\|_{\infty}=0.5$. Lines and dots indicate MSE mean averaged over 15 trials $\pm$ 95% CI.
(b) Parameter and test prediction MSE as a function of training time length $T$ and graph topology. The signal generated here was smooth (as per Equation 4.4), with $\|D\beta^{*}\|_{\infty}=0.5$. Lines and dots indicate MSE mean averaged over 15 trials $\pm$ 95% CI.
Figure 5: Parameter and test prediction MSE as a function of training time length $T$ for different graph topologies.
@@
We hold $T$ fixed (here, $T=10$) and increase the graph size ($m$). The results are shown in Figure 6 for a piecewise constant signal. We note that our TV estimator improves with larger $m$. By contrast, OLSind does not benefit from added nodes, while especially in the path or grid graph, the pooled OLS (OLSpooled) oversmoothes the signal, inducing a substantial parameter MSE compared to both graph-regularized penalties.
@@ 4.3 Real Data Experiments: EPA data
We evaluate the proposed estimator on data from the U.S. Environmental Protection Agency (EPA) national air-quality monitoring network. Specifically, we analyze daily arithmetic-mean concentrations of nitric oxide (NO), particulate matter with aerodynamic diameter $\leq 10\mu m$ (PM10), and daily mean temperature from 2010-2024.888Data are available from the EPA: https://www.epa.gov/outdoor-air-quality-data/download-daily-data. These measurements are obtained from 398 monitoring stations across the United States. We process the data by retaining measurements from primary instruments with at least 75% temporal coverage, removing observations flagged as special events, and applying variable-specific transformations to approximate normality (none for temperature, $\log$ for PM10, and a log–log transform $\log(\log(x+1))$ for NO). Stations are further filtered so that, within each year, fewer than 5% of observations are missing; this reduces the number of stations by roughly 45% on average (e.g., 221 stations in 2020; see Figure 7). Missing values are imputed via last observation carried forward. Using the reported geographic coordinates, we construct a spatial graph by connecting each station to its five nearest neighbors; the resulting graph is shown in Figure 7, with colors indicating NO on April 9, 2024 (day 100).
@@ Results.
Figure 10 illustrates the structural differences and implicit regularization of each method. With $T_{\text{train}}=15$ (a data-limited regime), the station-wise OLS fits are highly heterogeneous, as evidenced by the wide spread of coefficients. By contrast, pooled OLS yields a single global coefficient shared across stations. Both the graph total-variation (TV) denoiser and the Laplacian regularizer produce coefficients that vary smoothly over the spatial graph. As expected, TV recovers piecewise-constant (or piecewise-smooth) structure with sharp transitions aligned to graph boundaries, whereas Laplacian regularization yields a smoother, more diffuse spatial distribution.

### v2

@@ 2.3 Graph–geometry parameters
the combinatorial Laplacian and let $L^{\dagger}$ be the pseudoinverse of $L$. Denote the eigenvalues of $L$ by $\lambda_{1}(G)\geq\lambda_{2}(G)\geq\cdots\geq\lambda_{m-1}(G)>\lambda_{m}(G)(=0)$. Throughout, $\lambda_{m-1}(G)$ (or simply $\lambda_{m-1}$) stands for the algebraic connectivity—the smallest positive eigenvalue of $L$ (also known as the Fiedler eigenvalue of $G$). It is well known that $\lambda_{m-1}>0$ iff $G$ is connected, and more generally, $\lambda_{m-1}$ measures how well-knit the graph is. A small value indicates bottlenecks or sparse regions, making global information propagation harder.
@@ 2.4 Main results
which then allows us to suitably bound the terms $F_{1},F_{2}$ and $F_{3}$; see Appendix B for details.
@@
We now have the following useful corollary of Theorem 1 when each $A^{*}_{l}$ is stable. Its proof is outlined in Appendix B.
@@
    $\frac{1}{\sqrt{m}}\|\hat{a}-a^{*}\|_{2}\;\lesssim\;\frac{\mathfrak{M}}{\sqrt{mT}}\Big(1+\frac{\sqrt{s}}{\kappa_{S}}\Big),\qquad\Phi_{\mathcal{S}}=1+\frac{\sqrt{s}}{\kappa_{\mathcal{S}}}\;.$
@@
$T\gtrsim\max\left\{{\frac{\log^{2}(1/\delta)}{\Delta^{2}}\Phi_{\mathcal{S}}^{2}\log(dm/\delta),\frac{d^{2}+\log^{4}(1/\delta)}{m\,\Delta^{2}},\frac{\Phi_{\mathcal{S}}\,d\,\Delta_{G}}{\sqrt{m}}}\right\},$
where we recall $\Phi_{\mathcal{S}}=1+\|\widetilde{D}\,a^{*}\|_{1}$.
@@
$T\gtrsim\max\left\{{\frac{\log^{2}(1/\delta)}{\Delta^{2}}\Phi_{\mathcal{S}}^{2}\log^{2}(dm/\delta),\frac{d^{2}+\log^{4}(1/\delta)}{m\,\Delta^{2}},\Phi_{\mathcal{S}}\,d\,\Delta_{G}\sqrt{\frac{\log m}{m}}}\right\},$
where $\Phi_{\mathcal{S}}=1+\|\widetilde{D}\,a^{*}\|_{1}$. As before, consider for simplicity the very smooth regime where $\Phi_{\mathcal{S}},\Delta_{G}\leq c$ for a constant $c\geq 1$. Then if $\frac{m}{\log m}\gtrsim d^{2}+\log^{4}(1/\delta)$, we require $T\gtrsim\frac{\log^{2}(1/\delta)}{\Delta^{2}}\log^{2}(dm/\delta)$.
@@ 4 Experiments
In this section, we validate the performance of the graph total variation regularizer for estimating the LDS matrices $A^{*}_{l}$ on a series of synthetic experiments [^8] The code can be found at: https://github.com/donnate/LDS.. To this end, we consider linear dynamical systems on a graph on $m$ nodes where each node $l\in\{1,\dots,m\}$ is associated with a time series that evolves as
@@ 4.1 Experimental Protocol
We vectorize each $A^{*}_{l}\in\mathbb{R}^{d\times d}$ and stack them in a single vector $a^{*}=\big(\operatorname{vec}(A^{*}_{1})^{\top},\dots,\operatorname{vec}(A^{*}_{m})^{\top}\big)^{\top}\in\mathbb{R}^{md^{2}}$. Let $T_{\text{train}}$ denote the length of the time series used for training. For node $l$, with time series $X_{l}\in\mathbb{R}^{d\times T_{\text{train}}}$ and response $\widetilde{X}_{l}\in\mathbb{R}^{d\times T_{\text{train}}}$, the local design is $Q_{l}=(X_{l}^{\top}\otimes I_{d})\in\mathbb{R}^{(T_{\text{train}}d)\times d^{2}}$. We form $Q=\operatorname{blkdiag}(Q_{1},\dots,Q_{m})$ and $\tilde{x}=\operatorname{vec}(\widetilde{X}_{l})\in\mathbb{R}^{dT_{\text{train}}}.$
@@
Given a graph and a collection of LDS matrices $(A^{*}_{l})_{l=1}^{m}$, we simulate $T=T_{\text{train}}+2b+T_{\text{val}}+T_{\text{test}}$ steps per node from (4.1), with $b$ as a buffer parameter (here, chosen to be 100) that ensures that the validation and testing sets are (approximately) independent of one another and of the training set. We use the first $T_{\text{train}}$ steps for training, the next $T_{\text{val}}$ steps for validation, and the last $T_{\text{test}}$ for testing. Unless stated otherwise, we set $T_{\text{val}}=4$, $T_{\text{test}}=8$, and repeat over seeds $r=1,\dots,n_{\text{rep}}$.
@@ 4.2 Results
We also vary $T_{\text{train}}\in\{4,8,16,32,64,128,256\}$ (log-scale $x$-axis). We expect TV to clearly outperform OLSind at small $T_{\text{train}}$ by borrowing strength across nodes, while approaching OLSind as $T_{\text{train}}$ grows; OLSpooled is expected to perform well when $T_{\text{train}}$ is extremely small, as the benefits of the smoothing induced by fitting the estimator over the entire graph are expected to outweigh the bias incurred in this case. The results for the different topologies and signals are reported in Figures 5(a) and 5(b), and confirm our expected findings. In particular, for the path graph and dense Erdős–Rényi graph, the TV regularizer consistently outperforms all other estimators, including the Laplacian estimator. We also note that for the piecewise-constant signal, the graph TV estimator offers significant advantages in terms of estimation MSE over the Laplacian estimator (e.g. in the path and ER cases). Even in the 2D grid, as $T_{\text{train}}$ increases, we see that the TV regularizer improves upon the Laplacian. This is expected, as the TV regularizer’s ability to preserve sharp contrasts is advantageous for signals defined by a few sparse, high-magnitude differences rather than small, distributed ones.
(a) Parameter and test prediction MSE as a function of training time length $T_{\text{train}}$ and graph topology. The signal generated in these plots is piecewise constant (as per Equation 4.3), with $\|D\beta^{*}\|_{\infty}=0.5$. Lines and dots indicate MSE mean averaged over 15 trials $\pm$ 95% CI.
(b) Parameter and test prediction MSE as a function of training time length $T_{\text{train}}$ and graph topology. The signal generated here was smooth (as per Equation 4.4), with $\|D\beta^{*}\|_{\infty}=0.5$. Lines and dots indicate MSE mean averaged over 15 trials $\pm$ 95% CI.
Figure 5: Parameter and test prediction MSE as a function of training time length $T_{\text{train}}$ for different graph topologies.
@@
We hold $T_{\text{train}}$ fixed (here, $T_{\text{train}}=10$) and increase the graph size ($m$). The results are shown in Figure 6 for a piecewise constant signal. We note that our TV estimator improves with larger $m$. By contrast, OLSind does not benefit from added nodes, while especially in the path or grid graph, the pooled OLS (OLSpooled) oversmoothes the signal, inducing a substantial parameter MSE compared to both graph-regularized penalties.
@@ 4.3 Real Data Experiments: EPA data
We evaluate the proposed estimator on data from the U.S. Environmental Protection Agency (EPA) national air-quality monitoring network. Specifically, we analyze daily arithmetic-mean concentrations of nitric oxide (NO), particulate matter with aerodynamic diameter $\leq 10\mu m$ (PM10), and daily mean temperature from 2010-2024.999Data are available from the EPA: https://www.epa.gov/outdoor-air-quality-data/download-daily-data. These measurements are obtained from 398 monitoring stations across the United States. We process the data by retaining measurements from primary instruments with at least 75% temporal coverage, removing observations flagged as special events, and applying variable-specific transformations to approximate normality (none for temperature, $\log$ for PM10, and a log–log transform $\log(\log(x+1))$ for NO). Stations are further filtered so that, within each year, fewer than 5% of observations are missing; this reduces the number of stations by roughly 45% on average (e.g., 221 stations in 2020; see Figure 7). Missing values are imputed via last observation carried forward. Using the reported geographic coordinates, we construct a spatial graph by connecting each station to its five nearest neighbors; the resulting graph is shown in Figure 7, with colors indicating NO on April 9, 2024 (day 100).
@@ Results.
Figure 10 illustrates the structural differences and implicit regularization of each method. With $T_{\text{train}}=15$ (a data-limited regime), the station-wise OLS fits are highly heterogeneous, as evidenced by the wide spread of coefficients. By contrast, pooled OLS yields a single global coefficient shared across stations. Both the graph total-variation (TV) denoiser and the Laplacian regularizer produce coefficients that vary smoothly over the spatial graph. As expected, TV recovers piecewise-constant (or piecewise-smooth) structures with sharp transitions aligned to graph boundaries, whereas Laplacian regularization yields a smoother, more diffuse spatial distribution.